# Control of a commercial vehicle by a tetraplegic human using a bimanual brain-computer interface

### v1

@@ RESULTS
Figure 7: BCI driving scores and infractions. It started with an ideal 1.0 base score in each run, which was reduced each time an infraction was committed. (A and B) The teledriving scores and infractions of BCI-JJ using the unilateral cursor control for steering and speed of the commercial Ford Mustang Mach-E vehicle through four random routes in the Mcity facility encompassing an urban environment. (C and D) The simulated driving scores and infractions of BCI-JJ and the motor intact control group with the bimanual cursor-and-click control for steering, speed, and braking of a virtual vehicle on a town route modified from the CARLA Leaderboard 2.0. (E and F) The teledriving scores and infractions of BCI-JJ using bimanual control for steering, speed and braking of the Mach-E vehicle through an obstacle course. “CW” and “CCW” are short for clockwise and counterclockwise respectively.
BCI driving performance
Via the two reaction time tasks, we established that the reaction performance of BCI-JJ with index finger clicks decoded by our BCI system was at least as fast and reliable as the average performance of the motor intact participants clicking a computer mouse. In order to verify the robustness of our BCI system in dealing with more realistic applications, such as driving a vehicle, we outfitted a Ford Mustang Mach-E vehicle to be remotely controllable in real-world environments without traffic. Via teledriving, we first tested the speed and steering control with cursor movement decoded by our BCI system in a free-driving scenario through an urban test facility called Mcity (see Figure 3C). Then we increased the complexity of the BCI decoder from the unilateral cursor control for steering and speed to the bimanual cursor-and-click control, which enabled more precise full-stop braking via clicks with the left index finger effector and continuous steering and speed adjustments via the horizontal and vertical cursor movement, respectively, with the right thumb. To ensure the safety, precision, and complexity of our driving evaluations with bimanual control, we temporarily switched back to a laboratory setting and compared the driving proficiency of BCI-JJ to the motor intact control group using a virtual vehicle in a simulated town environment with busy traffic and a more sophisticated route modified from the CARLA Leaderboard 2.0 (see Figure 3E). Finally, we designed an obstacle course and applied teledriving with the same Mach-E vehicle which could verify our bimanual BCI driving system in a more applicable real-world setting mimicking many components of a standard driving test (see Figure 3G). The “BCI-enabled driving” section has a full description of the experimental settings for the three driving tasks.
We evaluated the following infractions, including the proportion of the completed route distance ($C$), the number of collisions ($N_{c}$), the number of lane deviations ($N_{l}$), and the number of running red traffic lights or stop signs if available ($N_{s}$). The infraction metric aggregated all these infractions using the following equation:
For both teledriving tasks in Mcity and through the obstacle course, BCI teledriving infractions were evaluated based on recorded videos by three independent evaluators from the California Institute of Technology, Blackrock Neurotech, and Ford Motor Company. Please note that if the vehicle was about to collide with some dangerous obstacle, $N_{c}$ would increase by 1, and the safety driver in the test vehicle would take over the vehicle control for a few seconds to move it back to the right track before the BCI teledriving could resume. If the vehicle was about to turn left at an intersection where a left-turn lane was available, it should merge to that left-turn lane before its left turn during BCI teledriving; otherwise, $N_{l}$ would increase by 1. If an evaluator was uncertain about the occurrence of any infraction component ($N_{c}$, $N_{l}$ or $N_{s}$) (e.g., whether there was a lane deviation at $t=45$ seconds in Run 2), he/she would mark it with a quantity increase of 0.5 instead of 1 (e.g., $N_{l}\mathrel{+}=0.5$).
@@
As a proof-of-concept, we started the evaluations of our BCI-enabled driving system with the unilateral cursor movement control for steering and speed of a commercial vehicle through a real-world urban environment without traffic. BCI-JJ, who was living in California, remotely controlled a Ford Mustang Mach-E vehicle through Mcity, a mock urban road environment located in Michigan. The goal of this Mcity teledriving task was to evaluate the continuousness and proficiency of BCI driving control in speed and steering, rather than to repeatedly evaluate reaction times or braking as we had done in the previous two reaction time tasks. Therefore, BCI-JJ was asked to ignore all traffic lights and stop signs in this urban test facility, focusing only on the unilateral cursor control for steering and speed with his right thumb. BCI-JJ completed four randomly mapped routes through Mcity, with one run per route across two sessions (see movies S1S2S3 and S4). Each route included various driving elements (e.g., turnings at intersections, roundabouts, U-turns, lane merges, etc., see Figure 3D and the “BCI-enabled driving” section for a full description).
With the BCI teledriving infractions in Mcity evaluated by the three independent evaluators, we applied the infraction metric to compute the cursor-only BCI teledriving score in Mcity using Eq. 1 in which $N_{s}$ was set to zero, because in this task no traffic light or stop sign was considered. Using BCI-decoded cursor movement for speed and steering control of the Mach-E vehicle, BCI-JJ completed each of the four random routes through Mcity in 2.5 to 7.5 minutes and obtained an average driving score of 0.78 (0.211 SD) across the four routes (see Figure 7A), with his driving performance improving over the course of each run (see Figure 7B).
@@
Using a right-tailed two sample t-test, we compared the composite simulated driving scores for BCI-JJ versus the motor intact control group. All participants applied the bimanual cursor-and-click control for steering, speed, and full-stop braking of the virtual vehicle. Across all runs, BCI-JJ, via our BCI control system, performed significantly better than the motor intact control group ($p<0.05$), with an average driving score of 0.924 (0.115 SD), compared to an average score of 0.823 (0.160 SD) for the motor intact control group, where a score of 1.0 is perfect (see Figure 7C, table S11, and movie S5 for a side-by-side comparison of the average simulated driving performance between BCI-JJ and a motor intact participant). Additionally, we used two sample t-tests to compare the number of average traffic violations between BCI-JJ and the motor intact control group. BCI-JJ had significantly fewer collisions per run, on average, than the control group ($p<0.05$). There were no significant differences between these two groups in the number of red light violations or lane deviations (see Figure 7D and table S11). These results reflect the previous findings of two reaction time tasks in which BCI-JJ reacted at least as fast and precisely as the motor intact control group. We also compared the simulated driving scores between BCI-JJ and all 20 motor intact participants individually, using a one-way ANOVA test followed by pairwise comparisons. We found BCI-JJ had his simulated town driving performance similar to 19 motor intact participants and only significantly better than 1 motor intact participant ($p<0.05$ Bonferroni corrected, see figure S6). These results indicate that the BCI control system can be used to control the essential functions of a vehicle as proficiently as motor intact people in a realistic driving environment.
Teledriving performance on an obstacle course
After establishing the safety and efficacy of our bimanual BCI control system in a complex simulated town driving task, we then demonstrated the capability of the bimanual BCI control system to navigate a commercial vehicle in a real-world driving scenario. To do this, BCI-JJ remotely controlled the same Mach-E vehicle through a loop-shaped driving obstacle course which included many components of a standard driving test (e.g., lane switches, turnings, full stops, roundabouts, obstacle avoidance, etc.) (see Figure 3H and the “BCI-enabled driving” section for a full description). In this obstacle-course teledriving task, BCI-JJ completed 9 runs per route for a total of four different route options (i.e., a clockwise or counterclockwise loop with or without a lane-switch segment) across three sessions (see movies S6S7S8 and S9 for one example run per route).
With the BCI teledriving infractions through the obstacle course evaluated by the three independent evaluators, we applied the infraction metric to compute the bimanual BCI teledriving score through the obstacle course using Eq. 1 in which $N_{s}$ represented the number of running stop signs. During all 36 runs through the obstacle course with the bimanual cursor-and-click control for steering, speed, and full-stop braking of the Mach-E vehicle, BCI-JJ received a composite driving score of 0.939 (0.0957 SD), where 1.0 is considered a perfect score (see Figure 7E). There were no significant differences in scores among the four different routes taken through the course (see Figure 7F), indicating that our bimanual BCI control system has the capability to remotely operate the basic controls of a commercial vehicle. To our knowledge, this is the first successful demonstration of full BCI control of a full-sized commercial vehicle in a real-world driving scenario.
@@ DISCUSSION
We conducted two reaction time tasks (i.e., the simple reaction time task and the braking reaction time task) by using different hand effectors of BCI-JJ with attempted clicks decoded by our BCI system and the right index finger of 20 motor intact participants with actual clicks via a computer mouse. Among tests of six hand effectors (i.e., right and left index fingers, ring fingers, and power grips) of BCI-JJ in the simple reaction time task, his right index finger had significantly faster reaction times and competitive performance measures. Compared with his left index finger in the braking reaction time task, BCI-JJ’s right index finger also had significantly faster braking reaction times; however, both index finger effectors had stable enough reaction performance in order to allow the vehicle’s braking in the face of emergencies. While the right thumb was selected for BCI-JJ to control the cursor movement for steering and speed changes in all three of our BCI driving tasks, his left index finger was selected to control the clicks for full-stop braking in both the simulated town driving task and the obstacle-course teledriving task. The differences in effector control are likely due to the placement of the arrays within MC, which has a rough somatotopic map for effectors, and degree of mixed selectivity of effectors, which is greater in PPC than MC [38, 39]. For fair comparisons, the same two effectors (i.e., the left index finger for clicks and the right thumb for cursor movement) were selected for motor intact participants during the simulated town driving task. Our reaction time task results have shown that the reaction times increase with task complexity for both BCI-JJ and the motor intact control group [40]. These results also indicate that reading human brain signals directly from the PPC and MC regions can provide reaction performance at least as fast and reliable as the motor intact behavior. This feature of our BCI system can be further applied to other time-sensitive devices.
Bimanual control with BCI
@@
Compared with largely focused single-effector control by many existing BCI systems [8, 41, 42], multi-effector movement allows greater functionality for people with paralysis. Our bimanually controlled BCI system can decode both the click movement via the left index finger and the cursor movement via the right thumb simultaneously and accurately using the partial least squares regression (PLSR) decoder with a feature extraction algorithm called FENet [34] (see the “BCI decoder system” section for a full description). We first documented the stability and accuracy of cursor and click control separately via two reaction time tasks and a free teledriving task in the Mcity closed test facility, respectively. Then we bimanually combined click control for braking and cursor control for speed and steering of a vehicle. In order to minimize the bias of comparisons between our BCI system and motor intact behavior for bimanual driving control, we asked all participants to navigate a virtual car through a simulated busy urban town. As judged by the simulator itself at the end of each run, the bimanual driving performance decoded by our BCI system from the PPC and MC regions of BCI-JJ turned out to be as stable and proficient as using a joystick by the motor intact control group. With the safety and efficacy of our bimanual BCI control established, we finally had a successful demonstration of the same BCI system teledriving a real-world vehicle through an obstacle course which contained many components of a standard driving test. Compared to existing bimanual control of BCI [10], the advance of our bimanual cursor-and-click BCI control system further enhances the degrees of freedom and complexities of real-world applications.
@@
Most existing intracortical BCI systems have been applied to controlled tasks within the laboratory or in simple home settings [1, 6, 8, 9, 33]. We show that current advances in intracortical BCIs extend the use of this technology to real-world scenarios, potentially improving the mobility of people with paralysis. Previous BCI systems for vehicle control suffered from temporal and spatial limitations [18, 22, 27, 31]. Our intracortical BCI system can provide proficient driving experiences for a tetraplegic participant not only in a simulated town environment with busy traffic but also in the real-world environments for the first time by remotely controlling a commercial vehicle. Using the bimanual cursor-and-click control via our BCI decoder, BCI-JJ has had his BCI-controlled driving performance in the CARLA 0.9.13 simulator comparable to 20 motor intact individuals with the same average age, according to the infraction components automatically recorded by the CARLA Leaderboard 2.0 at the end of each run. He has also performed BCI teledriving stably with a Ford Mach-E vehicle both in the Mcity closed test facility encompassing an urban environment without traffic (using the unilateral cursor control) and on an obstacle course (using the bimanual cursor-and-click control), according to the independent evaluations based on the recorded videos.
@@
Potential biases on the BCI versus motor intact control tests
Various factors can lead to biased comparisons, such as prior experience with driving simulators, joystick usage, and the level of involvement in the tasks between BCI-JJ and 20 motor intact participants. However, we made sure that the motor intact participants were selected with no gender bias and with their average age same as BCI-JJ’s. We only had the opportunity to test the BCI reaction time tasks and driving tasks with one tetraplegic participant who might already have faster reaction times before his injury than the average performance of the motor intact control group. BCI-JJ also had longer practice time than motor intact participants, although we tried to make sure the tasks were straightforward with a steep learning curve. We also do not have a comparison between our BCI system and the motor intact control group for driving a commercial vehicle in the real-world traffic environment.
@@
Our existing settings for both teledriving and simulated driving still had lags (e.g., feedforward control from the decoded brain signals or the joystick input to the vehicle, feedback control from the vehicle to the video display, etc.) and view restrictions (i.e., less than 180 degrees of the view ahead of the vehicle). As a result, we set the maximum speed of the simulated vehicle to 5 mph and the Ford Mach-E vehicle to 4 mph. Such low-speed vehicle control could ensure safety of the commercial vehicle and safety drivers, stable feedforward bimanual control in both a simulated busy town environment and a real-world obstacle course with narrow lanes, and smooth display of the video feedback. Completely eliminating all lags in the teleoperation pipeline might prove challenging, as this issue could arise at various stages, from data encoding and decoding to information transmission. Factors such as availability and bandwidth of the wireless network could also contribute to these lags. During BCI teledriving tests, the aggregated lags occasionally were non-negligible (i.e., greater than 1 to 2 seconds). When it happened, the safety driver immediately took control of the vehicle. Further development of a simulated driving setup with a VR system and/or an in-vehicle BCI setting for broader and more realistic views and shorter lags can help improve the driving realism and performance of both BCI and motor intact participants.
@@
Intracortical BCI control with high proficiency can improve safety and personal experiences when driving a semi-autonomous vehicle or using a robot [46]. By building on current results and improving multi-effector decoding, we can pursue more advanced engineering outcomes for driving with BCI. We can enrich our BCI system with a decoder of cognitive states which could provide a vigilance alert signal, trigger built-in vehicle safety features, and/or switch the vehicle control between autonomous driving and driver’s manual control. This advanced vehicle control system might eventually enable BCI participants to drive via their brain control inside the test vehicle instead of remotely with occasionally non-negligible telecommunication delay. In both simulated and physical urban scenarios with a semi-autonomous vehicle, the enhanced BCI system could continuously track the vigilance of each participant. If a participant is detected as having reduced attention, the system will alert them and strengthen the vehicle’s safety controls. Unlike an obstacle detection sensor (ODS) that comes with many commercial vehicles, such continuous BCI vigilance check can remind the driver to remain focused on the traffic at all times, rather than waiting until the vehicle gets too close to an obstacle or another vehicle. Although autonomous vehicles might be able to drive without human intervention in common scenarios, they might not respond adequately under special conditions (e.g., sudden appearances of other cars, pedestrians, obstacles, and unpredictable weather conditions). Therefore, the multi-effector BCI control and the autonomous driving features of the vehicle could be alternatively activated or used in combination under different situations. Such hierarchical BCI application can further improve the mobility of people with paralysis and advance the accuracy and flexibility of vehicle control.

### v2

@@ RESULTS
Figure 6: BCI driving scores and infractions. It started with an ideal 1.0 base score in each run, which was reduced each time an infraction was committed. (A and B) The teledriving scores and infractions of BCI-JJ using the single-effector cursor control for steering and speed of a Ford Mustang Mach-E vehicle through four random routes in the Mcity urban test facility without traffic. (C and D) The teledriving scores and infractions of BCI-JJ using the same cursor-movement control through an obstacle course. CW: clockwise. CCW: counterclockwise. (E and F) The simulated driving scores and infractions of BCI-JJ and the motor intact control group with the bimanual cursor-and-click control for steering, speed, and braking of a virtual vehicle in a busy town.
@@
As a proof-of-concept, we started the evaluations of our BCI-enabled driving system with the single-effector cursor control for steering and speed of a commercially available vehicle through a real-world urban environment without traffic. BCI-JJ, who was living in California, remotely controlled a Ford Mustang Mach-E vehicle through Mcity, a mock urban road environment located in Michigan. The goal of this Mcity teledriving task was to evaluate the continuousness and proficiency of BCI driving control in speed and steering, rather than to repeatedly evaluate reaction times or braking as we had done in the previous two reaction time tasks. Therefore, BCI-JJ was asked to ignore all traffic lights and stop signs in this urban test facility, focusing only on the cursor control for steering and speed with his right thumb. BCI-JJ completed four randomly mapped routes through Mcity, with one run per route across two sessions (see movies S1S2S3 and S4). Each route included various driving elements (e.g., turnings at intersections, roundabouts, U-turns, lane merges, etc., see Figure 3D and the “BCI-enabled driving” section for a full description).
With the BCI teledriving infractions in Mcity evaluated by the three independent human evaluators, we applied the infraction metric to compute the cursor-only BCI teledriving score in Mcity using Eq. 1 in which $N_{s}$ was set to zero, because in this task no traffic light or stop sign was considered. Using BCI-decoded cursor movement for speed and steering control of the Mach-E vehicle, BCI-JJ completed each of the four random routes through Mcity in 2.5 to 7.5 minutes and obtained an average driving score of 0.780 (0.211 SD) across the four routes (see Figure 6A), with his driving performance improving over the course of each run (see Figure 6B).
Teledriving performance on an obstacle course
After establishing the free-driving capability of our single-effector cursor-movement-based BCI control system for a commercially available vehicle in the Mcity closed test facility, we then demonstrated the capability of the same BCI control system to navigate the same vehicle on a more complex, predefined obstacle course. To do this, BCI-JJ remotely controlled the same Mach-E vehicle through a loop-shaped driving obstacle course which included many components of a standard driving test (e.g., lane switches, turnings, full stops, roundabouts, obstacle avoidance, etc.) (see Figure 3F and the “BCI-enabled driving” section for a full description). In this obstacle-course teledriving task, BCI-JJ completed 9 runs per route for a total of four different route options (i.e., a clockwise or counterclockwise loop with or without a lane-switch segment) across three sessions (see movies S5S6S7 and S8 for one example run per route).
With the BCI teledriving infractions through the obstacle course evaluated by the three independent evaluators, we applied the infraction metric to compute the BCI teledriving score through the obstacle course using Eq. 1 in which $N_{s}$ represented the number of running stop signs. During all 36 runs through the obstacle course with the single-effector cursor control for steering and speed of the Mach-E vehicle, BCI-JJ received a composite driving score of 0.939 (0.0957 SD), where 1.0 is considered a perfect score (see Figure 6C). There were no significant differences in scores among the four different routes taken through the course (see Figure 6D), indicating that our BCI control system has the capability to remotely operate the basic controls of a commercially available vehicle. To our knowledge, this is the first successful demonstration of full BCI control of a full-sized commercially available vehicle in a real-world driving scenario.
@@
Using a right-tailed two sample t-test, we compared the composite simulated driving scores for BCI-JJ versus the motor intact control group. All participants applied the bimanual cursor-and-click control for steering, speed, and full-stop braking of the virtual vehicle. Across all runs, BCI-JJ, via our BCI control system, performed significantly better than the motor intact control group ($p<0.05$), with an average driving score of 0.924 (0.115 SD), compared to an average score of 0.823 (0.160 SD) for the motor intact control group, where a score of 1.0 is perfect (see Figure 6E, table S11, and movie S9 for a side-by-side comparison of the average simulated driving performance between BCI-JJ and a motor intact participant). Additionally, we used two sample t-tests to compare the number of average traffic violations between BCI-JJ and the motor intact control group. BCI-JJ had significantly fewer collisions per run, on average, than the control group ($p<0.05$). There were no significant differences between these two groups in the number of red light violations or lane deviations (see Figure 6F and table S11). These results reflect the previous findings of two reaction time tasks in which BCI-JJ reacted at least as fast and precisely as the motor intact control group. We also compared the simulated driving scores between BCI-JJ and all 20 motor intact participants individually, using a one-way ANOVA test followed by pairwise comparisons. We found BCI-JJ had his simulated town driving performance similar to 19 motor intact participants and only significantly better than 1 motor intact participant ($p<0.05$ Bonferroni corrected, see figure S7). These results indicate that the BCI control system can be used to control the essential functions of a vehicle with the same proficiency level as motor intact people in a realistic driving environment.
A
B
C
D
E
Figure 7: Decoder stability and signal contributions. (A) Cross-validated $R^{2}$ values for the trained decoder on center-out data on each session of the Mcity teledriving task, the obstacle-course teledriving task, and the simulated town driving task. (B) Proportion of average Shapley values for BCI-JJ’s motor cortex (MC) and posterior parietal cortex (PPC), illustrating the contribution of each brain area to the decoder. (C) Normalized Shapley values for each channel across sessions, demonstrating the contribution of each channel to the decoder trained in each session. (D) Average normalized Shapley values for each channel plotted topographically for each electrode array implanted in BCI-$J$. (E) Average array impedance over years since implantation. Solid lines show the mean impedance of all channels for each electrode array across all sessions in which impedance was measured in a year; shaded areas represent one standard deviation (SD).
Decoder stability for driving
Each session began with a series of center-out tasks in which the computer progressively reduced its control over the cursor, ultimately giving control to the participant (see the “Cursor-and-click training” section). The $R^{2}$ of the trained decoders was stable across the three years of recordings (see Figure 7A) done for this study, with the linear regression not significantly different from a constant ($p=0.47$). We also calculated Shapley values for each driving session across the Mcity teledriving task, the obstacle-course teledriving task, and the simulated town driving task. Shapley values estimate each feature’s contribution to a decoder. By calculating the contributions of all channels for each implanted electrode (see Figure 7B), we found that MC contributed a mean of $64\%$ and PPC contributed a mean of $36\%$ ($0.8\%$ SD for both). Furthermore, the 32 channels that contributed the most to the decoder were the same across recorded sessions over the years of recordings (see Figure 7C). The locations of these high-contributing channels followed a clear topographic pattern in MC (see Figure 7D), but PPC did not show this topographic distribution.
Teledriving and simulated driving lags
In teleoperation tasks that require real-time feedback, a short latency between command generation and visual feedback is crucial for achieving precise control. To estimate the lag between BCI-JJ’s intentions and the video feedback of the vehicle’s movements, we measured the temporal difference between the overlay state and the vehicle’s movement direction as inferred from optic flow analysis of the videos on all three BCI-enabled driving tasks (see the “Lag estimation” section). This method accounted for delays caused by digital communication, video recording, and mechanical execution of commands.
A one-way ANOVA revealed a significant difference ($p<0.001$) in the estimated lags of the three tasks (see figure 9A). A Tukey’s HSD Test for multiple comparisons showed that the average lags among the three groups were significantly different. The Mcity teledriving task had the longest mean lag of -1333 ms (530 ms SD), the obstacle-course teledriving task had a mean lag of -750 ms (70 ms SD), and the simulated town driving task had the least mean lag of -200 ms (40 ms SD). We also found a strong negative Pearson correlation coefficient of -0.953 ($p<0.001$) between the magnitude of the lags and the scores adjusted for the route lengths (see the “Lag estimation” section).
@@ DISCUSSION
We conducted two reaction time tasks (i.e., the simple reaction time task and the braking reaction time task) by using different hand effectors of BCI-JJ with attempted clicks decoded by our BCI system and the right index finger of 20 motor intact participants with actual clicks via a computer mouse. Among tests of six hand effectors (i.e., right and left index fingers, ring fingers, and power grips) of BCI-JJ in the simple reaction time task, his right index finger had significantly faster reaction times and competitive performance measures. Compared with his left index finger in the braking reaction time task, BCI-JJ’s right index finger also had significantly faster braking reaction times; however, both index finger effectors had stable enough reaction performance in order to allow the vehicle’s braking in the face of emergencies. While the right thumb was selected for BCI-JJ to control the cursor movement for steering and speed changes among all three of our BCI driving tasks, his left index finger was selected to control the clicks for full-stop braking in the simulated town driving task. The differences in effector control are likely due to the placement of the arrays within MC, which has a rough somatotopic map for effectors, and degree of mixed selectivity of effectors, which is greater in PPC than MC [38, 39]. For fair comparisons, the same two effectors (i.e., the left index finger for clicks and the right thumb for cursor movement) were selected for motor intact participants during the simulated town driving task. Our reaction time task results have shown that the reaction times increase with task complexity for both BCI-JJ and the motor intact control group [40]. These results also indicate that reading human brain signals directly from the PPC and MC regions can provide reaction performance at least as fast and reliable as the motor intact behavior. This feature of our BCI system can be further applied to other time-sensitive devices.
Bimanual cursor-and-click control with BCI
@@
Compared with largely focused single-effector control by many existing BCI systems [8, 41, 42], multi-effector movement allows greater functionality for people with paralysis. Other groups have done bimanual BCI studies in the lab such as a delayed-movement task [10] or virtual quadcopter control to reach specific points without a designated route, any other vehicles or any real-world traffic elements [43], whereas our bimanually controlled BCI system has been successfully applied to driving a virtual vehicle to follow a designated route through heavy traffic in a simulated town environment with our BCI performance comparable to a motor intact control group. Our bimanual control involves both the click movement via the left index finger and the cursor movement via the right thumb simultaneously and accurately using the partial least squares regression (PLSR) decoder with a feature extraction algorithm called FENet [34] (see the “Feature extraction” section for a full description). We first documented the stability and accuracy of cursor and click control separately via two reaction time tasks and two teledriving tasks with a commercially available vehicle, respectively. The first teledriving task occurred in the Mcity closed test facility, whereas the second teledriving task took place on an obstacle course which contained many components of a standard driving test. Then we bimanually combined click control for braking and cursor control for speed and steering of a vehicle. In order to minimize the bias of comparisons between our BCI system and motor intact behavior for bimanual driving control, we asked all participants to navigate a virtual agent vehicle through a simulated busy urban town. As judged by the simulator itself at the end of each run, the bimanual driving performance decoded by our BCI system from the PPC and MC regions of BCI-JJ turned out to have the same level of stability and proficiency as using a joystick by the motor intact control group. We achieved stable bimanual decoding, performing in trials that lasted nearly two orders of magnitude longer than typical cursor control tests, which has not been demonstrated before. Beyond this, most BCI experiments report results over a few days or weeks, whereas here we show that our system’s performance remained stable for several years.
@@
Most existing intracortical BCI systems have been applied to controlled tasks within the laboratory or home settings [1, 6, 8, 9, 33]. We show that current advances in intracortical BCIs extend the use of this technology to real-world scenarios, potentially improving the mobility of people with paralysis. Previous BCI systems for vehicle control suffered from temporal and spatial limitations [18, 22, 27, 31]. We developed a modular system that included neural signal processing, bimanual cursor-and-click decoders, a graphical user interface, communication protocols, and engineering modules, enabling vehicle control. Our intracortical BCI system can provide first-of-its-kind driving experiences for a tetraplegic participant in the real-world environments by remotely controlling the steering and speed of a commercially available vehicle. BCI-JJ has performed single-effector cursor-movement-based BCI teledriving stably with a Ford Mach-E vehicle both in the Mcity closed test facility encompassing an urban environment via randomly mapped routes and on a predefined obstacle course including many components of a standard driving test, according to the independent human evaluations based on the recorded videos. Our BCI system also allows BCI-JJ to drive in a simulated town environment with heavy traffic by controlling the speed, steering, and full-stop braking of a virtual vehicle. Using the more advanced bimanual cursor-and-click control via our BCI decoder, BCI-JJ has had his BCI-controlled driving performance in the simulated town via CARLA Leaderboard 2.0 comparable to the joystick-controlled performance of 20 motor intact individuals with their average age the same as BCI-JJ’s, according to the infraction components automatically recorded by the CARLA Leaderboard 2.0 at the end of each run. It is also important to note that the NeuroPort Electrodes had been implanted in BCI-JJ for three years when this study started, whereas the performance of our decoder remained stable across three years of recordings for this study, contributing to reliable translatability from neuroscience to BCI real-world applications in the long term.
@@
Driving test settings
The teledriving test settings were less practical than the driving experiences that would normally occur in our daily lives. For the Mcity teledriving task, we asked BCI-JJ to complete teledriving of the Ford Mustang Mach-E vehicle on randomly mapped routes, meaning he could ignore all the traffic lights and stop signs in this closed test facility and instead focus on keeping the vehicle in the lane and avoiding collisions. For the obstacle-course teledriving task, we added stop signs and narrower lanes for the same vehicle to follow. To evaluate whether our teledriving setup served as a proof-of-concept for our BCI system’s ability to reliably control the movement of a commercially available vehicle in real-world scenarios, we used domain-specific metrics designed to assess driving performance rather than general BCI metrics. This includes the CARLA Leaderboard 2.0 infraction measures and relevant parts of the California DMV driving test checklist [44] (see the “Safety criteria for driving tasks” section).
CARLA Leaderboard 2.0 was designed for autonomous driving tests, incorporating various in-vehicle sensors and artificial intelligence algorithms. Limitations of the simulation environment prevented either BCI-JJ or the motor intact control group from achieving perfect driving performance in the simulated town driving task with bimanual cursor-and-click control on the overlay. Some of the limitations are intrinsic to the CARLA simulator, such as a restricted and fixed field of view for the virtual agent vehicle, no yielding rule for other autonomous vehicles, and partial incoherence for the physics between the virtual agent vehicle and a physical vehicle. Moreover, some limitations are related to the experimental setup, such as the joystick controller and the monitor. However, the simulator was a valuable testing framework, as it provided objective evaluations of many practical features of real-world driving environments, including heavy traffic, functioning traffic lights, and computer-measurable infraction metrics.
Potential biases on the BCI versus motor intact control comparisons
Various factors can lead to biased comparisons, such as prior experience with driving simulators, joystick usage, and the level of involvement in the tasks between BCI-JJ and 20 motor intact participants. However, we made sure that the motor intact participants were selected with no gender bias and with their average age the same as BCI-JJ’s. We only had the opportunity to test the BCI reaction time tasks and driving tasks with one tetraplegic participant who might already have faster reaction times before his injury than the average performance of the motor intact control group. BCI-JJ also had longer practice time than motor intact participants, although we tried to make sure the tasks were straightforward with a steep learning curve. We also do not have a comparison between our BCI system and the motor intact control group for driving a commercially available vehicle in the real-world traffic environment.
@@
We found that the teledriving system we used experienced significant lags during both the Mcity and the obstacle-course tasks. Humans are capable of adjusting their motor actions to visual feedback lags greater than 200 ms, but longer lags tend to increase the perceived complexity of the task [45, 46]. These long lags between motor intention and visual feedback during the teledriving tasks made driving the vehicle more challenging, as the participant did not receive timely visual information on the effect of his actions. This delayed sensorimotor integration can result in slower motor responses and increased overcorrections, negatively affecting control precision.
As a result, we set the maximum speed of the simulated vehicle to 5 mph and the Ford Mach-E vehicle to 4 mph. Such low-speed vehicle control could ensure safety of the commercially available vehicle and safety drivers, stable control in both a simulated busy town environment and a real-world obstacle course with narrow lanes, and smooth display of the video feedback. Completely eliminating all lags in the teleoperation pipeline might prove challenging, as this issue was mainly affected by digital communication delays. Factors such as availability and bandwidth of the wireless network contributed to these lags. Further development of a simulated driving setup with a VR system and/or an in-vehicle BCI driver setting for broader and more realistic views and shorter lags can help improve the driving realism and performance of both BCI and motor intact participants.
Long-term recording quality and decoder stability
We found that the performance of our decoder was stable over three years of this study (see Figure 7A). It is important to note that BCI-JJ had been implanted with the NeuroPort Electrodes for three years by the time the study started, with the impedance on the channels diminishing at a much slower rate than during the initial years after implantation (see Figure 7E). Even with the downgraded recording signal quality, the high stability of the decoder could still benefit our real-world BCI applications.
@@
In-vehicle driving with multi-effector BCI control
We applied the cursor-movement-based BCI control with the right thumb effector for steering and speed adjustments when teledriving a commercially available vehicle, and then added the additional click control with the left index finger for full-stop braking in the simulated town driving task. The bimanual setting has been proved to allow BCI-JJ to drive with the same proficiency level as the motor intact control group in a virtual town with heavy traffic, so we would like to apply the same bimanual setting or even more advanced multi-effector control to driving a commercially available vehicle in the future. The lags between between BCI-JJ’s intentions and the video feedback of the vehicle’s movements in the three driving tasks have a negative correlation with the adjusted driving scores based on the route length. Instead of just remotely driving a commercially available vehicle, we would like to further test our BCI control system with a BCI participant sitting and driving inside the vehicle for enhanced mobility, wider field of view, and reduced lags.
@@
Intracortical BCI control with high proficiency can improve safety and personal experiences when driving a semi-autonomous vehicle or using a robot [50]. By building on current results and improving multi-effector decoding, we can pursue more advanced engineering outcomes for driving with BCI. We can enrich our BCI system with a decoder of cognitive states which could provide a vigilance alert signal, trigger built-in vehicle safety features, and/or switch the vehicle control between autonomous driving and driver’s manual control. This advanced vehicle control system might eventually enable BCI participants to drive via their brain control inside the test vehicle instead of remotely with occasionally non-negligible telecommunication delay. In both simulated and physical urban scenarios with a semi-autonomous vehicle, the enhanced BCI system could continuously track the vigilance of each participant. If a participant is detected as having reduced attention, the system will alert them and strengthen the vehicle’s safety controls. Unlike an obstacle detection sensor (ODS) that comes with many commercially available vehicles, such continuous BCI vigilance check can remind the driver to remain focused on the traffic at all times, rather than waiting until the vehicle gets too close to an obstacle or another vehicle. Although autonomous vehicles might be able to drive without human intervention in common scenarios, they might not respond adequately under special conditions (e.g., sudden appearances of other cars, pedestrians, obstacles, and unpredictable weather conditions). Therefore, the multi-effector BCI control and the autonomous driving features of the vehicle could be alternatively activated or used in combination under different situations. Such hierarchical BCI application can further improve the mobility of people with paralysis and advance the accuracy and flexibility of vehicle control.